# UrbanWorld: An Urban World Model for 3D City Generation

### v1

@@ Abstract
Cities, as the most fundamental environment of human life, encompass diverse physical elements such as buildings, roads and vegetation with complex interconnection. Crafting realistic, interactive 3D urban environments plays a crucial role in constructing AI agents capable of perceiving, decision-making, and acting like humans in real-world environments. However, creating high-fidelity 3D urban environments usually entails extensive manual labor from designers, involving intricate detailing and accurate representation of complex urban features. Therefore, how to accomplish this in an automatical way remains a longstanding challenge. Toward this problem, we propose UrbanWorld, the first generative urban world model that can automatically create a customized, realistic and interactive 3D urban world with flexible control conditions. UrbanWorld incorporates four key stages in the automatical crafting pipeline: 3D layout generation from openly accessible OSM data, urban scene planning and designing with a powerful urban multimodal large language model (Urban MLLM), controllable urban asset rendering with advanced 3D diffusion techniques, and finally the MLLM-assisted scene refinement. The crafted high-fidelity 3D urban environments enable realistic feedback and interactions for general AI and machine perceptual systems in simulations. We are working on contributing UrbanWorld as an open-source and versatile platform for evaluating and improving AI abilities in perception, decision-making, and interaction in realistic urban environments.
@@ 1 Introduction
Cities are the most complex human-centric environments, characterized by their intricate structures, diverse elements, and dynamic interactions. Creating near-realistic 3D urban world environments is a fundamental and pivotal technique for broad research and real applications across various domains such as AI agents (Yang et al., 2024), urban planning (Schrotter & Hürzeler, 2020), urban simulation (Xu et al., 2023) and metaverse (Allam et al., 2022). Traditionally, achieving this involves expensive labor costs for human designers on detailed asset modeling, texture mapping, and scene composition. With the advancement of generative AI, there have emerged more automatic approaches for 3D scene generation based on volumetric rendering(Lin et al., 2023; Xie et al., 2024) and diffusion models(Deng et al., 2023; Lu et al., 2024). These approaches have revolutionized the paradigm of 3D scene generation, alleviating the high costs of manual design. However, the crafted 3D scenes are only visually appealing videos, which are significantly different from the real embodied physical world. Regarding this issue, a recent series of methods known as world models have emerged, preliminarily focusing on autonomous driving scenes (Hu et al., 2023; Wang et al., 2023). These models are shown to possess the capability of understanding the scene dynamics and predicting future states, uplifting the interactivity of 3D scene generation. However, there is still a large gulf between the created urban environments and the real urban world in which humans live. To sum up, there is still a long way from the actual “urban world models”, which we define as models able to create urban environments that are (1) realistic and interactive (2) customizable and controllable (3) capable of supporting embodied agent learning.
@@
Urban world models are of great significance in developing embodied intelligence and Artificial General Intelligence (AGI). Firstly, it is promising to bridge the gap between virtual environments and the real world, enabling embodied agents to interact with and learn from richly detailed, realistic urban environments. Secondly, by crafting synthetic 3D urban environments, researchers can gain complete control over data generation, with full access to all generative parameters. Machine perceptual systems can thus be trained on tasks that are not well suited to conduct in the real world or require various environments. Finally, a sophisticated urban world model can simulate a wide variety of environments, from bustling city centers to quiet residential neighborhoods, with realistic visual appearances of physical infrastructures such as buildings, roads, and natural spaces. This is crucial to avoid overfitting and creating agents with high generalization in diverse and dynamic environments. Some preliminary explorations have been witnessed in some commercial platforms like Omniverse [^1] https://www.nvidia.com/en-us/omniverse/, an open platform for creating and simulating detailed 3D world environments. However, there are no specialized urban world models for crafting interactive 3D urban environments, hindering the development of agents in the realistic urban environments.
@@
Toward this issue, we propose UrbanWorld, a generative urban world model that can automatically create realistic, controllable and embodied 3D urban environments just from a text and OpenStreetMap [^2] https://www.openstreetmap.org/ (OSM) prompt. In detail, there are four key modules in the framework of UrbanWorld. Firstly, UrbanWorld automatically generates 3D urban layouts and conducts detailed asset processing based on open-accessible OSM data via Blender [^3] https://www.blender.org/. Then, UrbanWorld adopts a fine-tuned urban-specific multimodal large language model (called Urban MLLM) to effectively plan and design urban scenes following user instructions, generating detailed textual descriptions of urban elements. Next, UrbanWorld integrates a 3D asset renderer based on texture diffusion and refinement, flexibly controlled by textual and visual conditions. Finally, to further optimize the visual appearance, UrbanWorld utilizes Urban MLLM to scrutinize the crafted 3D urban scenes, generating detailed suggestions for refinement and activate an additional iteration of rendering. This framework fully exploits the controllable generation capabilities of diffusion models and the reasoning and planning abilities of MLLMs, contributing to the high-fidelity urban environment rendering and superior generation flexibility.
The created urban environment features detailed and realistic visual representations of urban infrastructures, including buildings, roads, and natural areas, greatly benefiting the deployment of AI systems. Notably, it can support near-realistic physical interactions between urban environments and agents. It is tailored to enhance a range of embodied abilities including sensory grounding, perception and decision-making. For example, embodied agents can be deployed and trained for object recognition, route planning and automatic navigation via interaction with the realistic urban environment. Besides, UrbanWorld can serve as an open platform to support the creation and manipulation of more advanced 3D urban environments, facilitating the advancement of broad AI communities.
@@
  • We provide UrbanWorld as an open-source tool and contribute a 3D asset dataset with various urban environments, which facilitates the advancement of broad AI-related research, such as embodied intelligence and AI agents, laying the groundwork for advancing AGI.
@@ 2 Related Works
Table 1: Comparison between existing works for 3D urban scene generation and UrbanWorld from four aspects: text-controllable, image-controllable, creative (i.e., whether new assets can be created) and embodied/interactive (whether the created urban environment is physically interactive).
Method Text-controllable Image-controllable Creative Embodied/Interactive
SceneDreamer (Chen et al., 2023) $\times$ $\checkmark$ $\checkmark$ $\times$
PersistentNature (Chai et al., 2023) $\times$ $\checkmark$ $\checkmark$ $\times$
Infinicity Lin et al. (2023) $\times$ $\checkmark$ $\checkmark$ $\times$
CityDreamer (Xie et al., 2024) $\times$ $\times$ $\checkmark$ $\times$
CityGen (Deng et al., 2023) $\times$ $\times$ $\checkmark$ $\times$
SceneCraft (Hu et al., 2024) $\checkmark$ $\times$ $\times$ $\checkmark$
CityCraft Deng et al. (2024) $\checkmark$ $\times$ $\times$ $\checkmark$
UrbanWorld $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
3D urban scene generation aims to create realistic 3D urban environments with sophisticated urban planning and visual element design, usually requiring high human efforts such as complex asset modeling, texture mapping, and scene composition. With the advancement of deep learning techniques, recently there are three lines of work trying to achieve this in an automated way, including NeRF-based methods (Lin et al., 2023; Xie et al., 2024; Chen et al., 2023), diffusion-based methods (Deng et al., 2023; Inoue et al., 2023; Wu et al., 2024) and professional software script-based methods (Zhou et al., 2024; Hu et al., 2024). An overview of the method comparison is shown in Table 1. NeRF-based methods implicitly represent the urban scene and perform the volumetric rendering for the neural fields. For example, CityDreamer (Xie et al., 2024) first separates the scene into buildings and backgrounds then introduces different types of neural fields for asset rendering. These methods can produce a high-quality visual appearance while potentially losing geometric fidelity. Diffusion-based methods utilize diffusion models to generate city layouts or urban scenes. CityGen (Deng et al., 2023) provides an end-to-end pipeline to create diverse 3D city layouts with Stable Diffusion. These methods are creative in generating scene images or videos, but hard to obtain embodied 3D environments, limiting the practical usages. Recently, some professional software script-based methods have been proposed, trying to develop an automatic agentic workflow using LLMs to control the professional software for scene creation. A representative example is Scenecraft (Hu et al., 2024), which establishes an LLM-based agent to translate textual descriptions into Python scripts for scene creation in Blender. A more recent work CityCraft (Deng et al., 2024) adopts LLMs in designing and organizing 3D urban environments from off-the-shelf asset libraries. Such approaches are effective but only create urban environments by retrieving and organizing existing assets, unable to flexibly create new assets when necessary. Differently, our approach adopts and integrates both diffusion-based and professional software script-based methods, providing an effective, controllable and creative way for 3D urban world generation.
@@
A persistent objective in AI research has been to develop machine agents capable of engaging with various environments in 3D space like humans. Toward this goal, researchers have been devoted to building various interactive world simulators in the format of videos (Bruce et al., 2024) or embodied environment (Shen et al., 2022). Existing world simulation environments and platforms are mostly for indoor scenes (Puig et al., 2018; Xia et al., 2018; Kolve et al., 2017; Xia et al., 2020). Differently, Threedworld (Gan et al., 2020) pays attention to creating outdoor environments by retrieving and compositing objects from an existing asset library. When it comes to urban scenes which is the most important open environment, existing works mostly focus on the generative world model for autonomous driving capable of learning scene dynamics and understanding the geometry of the physical world (Hu et al., 2023; Wang et al., 2023; 2024). However, these models can only generate new scenes in the format of videos, hard to provide an embodied and interactive urban environment for real use. UGI (Xu et al., 2023) takes a forward step toward urban world simulation, proposing to develop an embodied urban environment for agent development. Although it conceptualized some relevant ideas, there is a lack of specific frameworks for implementation. To address this challenge, we propose UrbanWorld, which is expected to facilitate the construction of diverse embodied urban environments with controllable and refined visual appearance, supporting agent development or simulation in various urban scenes.
@@ 3 Methodology
Figure 1: Illustration of the whole framework of UrbanWorld, including four key components: (A) OSM-guided urban layout generation; (B) MLLM-empowered urban scene design; (C) Diffusion-based urban asset texture rendering; (D) MLLM-assisted scene refinement.
@@
There are three main challenges to solve for building an “urban world model”: efficient embodied environment construction, professional urban scene planning and high-quality texture generation. Towards these objectives, UrbanWorld introduces four key components: (1) OSM-guided urban layout generation, an automatic 2D-to-3D transformation module based on globally open-accessible OSM data, which can address the first challenge. (2) MLLM-empowered urban scene design, which exploits the superior urban scene understanding ability of the trained urban MLLM to draft reasonable urban scenes like human designers towards the second challenge. (3) Controllable diffusion-based urban asset texture renderer, a flexible urban asset rendering based on 3D diffusion following customized prompts. (4) MLLM-assisted urban scene refinement, a final reflection module to further improve the scene design, inspired by the iterative revision in the standard operation process of human designers. The last two components contribute to high-fidelity textures of 3D assets, effectively tackling the third challenge. The overview of UrbanWorld is illustrated in Figure 1.
3.1 OSM-guided Urban Layout Generation
@@
Considering the easy accessibility and global coverage of OSM data, UrbanWorld is mainly developed based on OSM to generate 3D urban layouts. OSM data contains a wealth of information, mainly including the geographic locations and attributes of roads, buildings, vegetation, forests, water, and other infrastructure elements. The contained urban assets such as buildings, forest, vegetation, water and roads are then separated as independent objects for subsequent element-wise rendering. In this step, UrbanWorld also records the object center location for further reorganization of assets, making it match the real urban layout.
3.2 MLLM-empowered Urban Scene Design
@@
Aiming to effectively craft customized urban environments, UrbanWorld integrates an advanced urban MLLM trained on extensive urban street-view imagery data. Specifically, we first collected urban street-view images globally and labeled the corresponding textual descriptions with GPT-4 and then conducted a manual check and filtered the low-quality data. Then we fine-tune an open-source MLLM, LLaVA-1.5 (Liu et al., 2024) on around 100K image-text pairs from the collected dataset. We have validated that the obtained urban MLLM possesses outstanding performance on urban scene understanding tasks such as image captioning and scene classification thus can benefit the urban scene analysis and design. In UrbanWorld, the urban MLLM is introduced to act as a human-like designer, which automatically drafts high-quality and detailed urban scene descriptions, ensuring the urban scenes are visually coherent. Specifically, taking a simple textual instruction (e.g., a teaching area in the university) from users and the selected OSM layout image as input, UrbanWorld calls the urban MLLM with carefully designed prompts and returns diverse detailed descriptions about visual appearance and materials for each asset. The produced asset descriptions will be used as the condition controlling the later rendering process.
@@ 3.3 Controllable diffusion-based urban asset texture rendering
Rendering a large-scale urban scene is challenging due to the existence of complex elements and relations, whereby the scene-level rendering will inevitably result in mismatching and low-resolution texture. Therefore, we follow the element-wise rendering principle to ensure the rendering quality. Simultaneously to speed up the rendering process, we merge some urban element types and finally define four main categories: buildings, roads and paths, forest and vegetation, and water. We implement the rendering with a controllable diffusion-based method consisting of two stages: UV texture generation and texture refinement detailed as shown in Figure 2.
Following the previous procedures, we have obtained the untextured 3D mesh of the whole urban scene $S$ and each asset $S_{i}$. For each type of element, we now have corresponding textual descriptions $t_{i}$ from the MLLM. here the model also supports taking reference image $r_{i}$ as prompts to control the generation. The UV map of the i-th asset $S_{i}$ is denoted as $U_{i}$.
We first set a series of camera views $v_{i}$ = $\{v_{k}\}_{k=1}^{N}$ to capture the multi-view appearance of the object. Next, we utilize the depth-aware ControlNet (Zhang et al., 2023) to control a 2D diffusion model $F$ to generate an image $I_{i}$ showing the visual appearance on different views, together with the condition $c_{i}\in\{t_{i},r_{i},(t_{i},r_{i})\}$:
@@
where $z$ is the latent embedding for the diffusion process, the depth map from different views $d_{i}$ is obtained from the rendering process $d_{i}=P(S_{i},v_{i})$. Then we conduct a reverse process $P^{-1}$ of rendering to back-project $I_{i}$ into the UV texture space from each view:
$U_{i}=P^{-1}(v_{i};I_{i};S_{i}),$ (2)
Then we crop $U_{i}$ into texture maps from each view $\{U_{i}^{k}\}_{k=1}^{N}$, each containing unique textual information from different views. Subsequently, we merge these texture maps into a single texture map $U_{i}$:
@@
Up to now, we have obtained the preliminary texture map for the asset $S_{i}$, while practically we have found that there are still some untextured areas on the object due to the discrete sampling of views, especially for buildings with many faces. Inspired by the inpainting capability of diffusion models, we introduce an additional UV texture inpainting process to get complete and natural textures. However, such inpainting can not be directly achieved with general diffusion-based inpainting considering there are strict position mapping requirements for UV texture maps. Therefore, we follow the UV inpainting process of general 3D objects  (Zeng et al., 2024), and introduce a position map-guided building UV inpainting process.
@@
To be specific, we add a UV position map encoder $E_{V}$ to encode the position map $V_{i}\in\mathbb{R}^{H\times W\times 3}$, indicating the adjacency relation of the UV texture fragments, where $E_{V}$ shares the same architecture of the image encoder in the diffusion model. Then we curate a set of paired UV position maps and UV texture maps for urban assets with complex surfaces and train the position encoder following the pipeline of ControlNet (Zhang et al., 2023). With the control of UV position maps, it’s expected to achieve accurate and reasonable inpainting for UV texture maps. The UV inpainting process is formulated as follows:
$U_{i}^{refine}=F(c_{i};U_{i};E_{V}(V_{i})).$ (4)
With the above texture generation and completion steps, UrbanWorld can produce coherent and high-fidelity textures for various urban elements.
For a better visual aesthetic of the rendering, we conduct image upscaling with ControlNet-tile to further enhance the structure sharpness and realism of the texture map, producing more detailed and realistic appearances for urban assets.

### v2

@@ Abstract
Cities, as the essential environment of human life, encompass diverse physical elements such as buildings, roads and vegetation, which continuously interact with dynamic entities like people and vehicles. Crafting realistic, interactive 3D urban environments is essential for nurturing AGI systems and constructing AI agents capable of perceiving, decision-making, and acting like humans in real-world environments. However, creating high-fidelity 3D urban environments usually entails extensive manual labor from designers, involving intricate detailing and representation of complex urban elements. Therefore, accomplishing this automatically remains a longstanding challenge. Toward this problem, we propose UrbanWorld, the first generative urban world model that can automatically create a customized, realistic and interactive 3D urban world with flexible control conditions. Specifically, we design a progressive diffusion-based rendering method to produce 3D urban assets with high-quality textures. Moreover, we propose a specialized urban multimodal large language model (Urban MLLM) trained on realistic street-view image-text corpus to supervise and guide the generation process. UrbanWorld incorporates four key stages in the generation pipeline: flexible 3D layout generation from OSM data or urban layout with semantic and height maps, urban scene design with Urban MLLM, controllable urban asset rendering via progressive 3D diffusion, and MLLM-assisted scene refinement. We conduct extensive quantitative analysis on five visual metrics, demonstrating that UrbanWorld achieves state-of-the-art generation realism. Next, we provide qualitative results about the controllable generation capabilities of UrbanWorld using both textual and image-based prompts. Lastly, we verify the interactive nature of these environments by showcasing the agent perception and navigation within the created environments. We contribute UrbanWorld as an open-source tool available at https://github.com/Urban-World/UrbanWorld.
@@ 1 Introduction
Cities are the most complex human-centric environments, characterized by their intricate spatial structures, heterogeneous components such as buildings, infrastructure, and public spaces, and dynamic interactions between these components and human activities. Creating near-realistic 3D urban world environments is a fundamental technique for broad research and real applications across various domains such as Artificial General Intelligence [1], AI agents [2], embodied AI [3], urban simulation [4] and metaverse [5]. Traditionally, achieving this involves expensive labor costs for human designers on detailed asset modeling, texture mapping, and scene composition. With the advancement of generative AI, there have emerged some automatic approaches for 3D scene generation based on volumetric rendering[6, 7] and diffusion models[8, 9]. These approaches have revolutionized the paradigm of 3D scene generation, alleviating the high costs of manual design. However, the crafted 3D scenes are limited to the video format and unable to provide embodied and interactive environments. Regarding this issue, a recent series of methods known as world models have emerged, preliminarily focusing on autonomous driving scenes [10, 11]. These models are shown to possess the capability of understanding the scene dynamics and predicting future states, uplifting the interactivity of 3D scene generation. However, there is still a large gap between the created urban environments and the real urban world in which humans live. To sum up, there is still a long way from the actual “urban world models”, which we define as models able to create urban environments that are (1) realistic and interactive (2) customizable and controllable (3) capable of supporting embodied agent learning. Table 1 provides a comprehensive review of existing 3D city generation methods from these key perspectives, highlighting that no existing approach can fully satisfy these requirements.
@@
Urban world models are of great significance in developing embodied intelligence and Artificial General Intelligence (AGI). Firstly, it is promising to bridge the gap between virtual environments and the real world, enabling embodied agents to interact with and learn from richly detailed, realistic urban environments. Secondly, by crafting synthetic 3D urban environments, researchers can gain complete control over data generation, with full access to all generative parameters. Machine perceptual systems can thus be trained on tasks that are not well suited to conduct in the real world or require various environments. Finally, a sophisticated urban world model can simulate a wide variety of environments, from bustling city centers to quiet residential neighborhoods, with realistic visual appearances of physical infrastructures such as buildings, roads, and natural spaces. This is crucial to avoid overfitting and creating agents with high generalization in diverse and dynamic environments. However, there are no specialized urban world models for automatically crafting interactive 3D urban environments, hindering the advancement of AI abilities toward general intelligence through embodied learning in complex open-world environments.
@@
Toward this issue, we propose UrbanWorld, a generative urban world model that can automatically create realistic, controllable and embodied 3D urban environments from user instructions and urban layout data in various format such as OpenStreetMaphttps://www.openstreetmap.org/ (OSM) and layouts with semantic and depth maps  [8, 12]. In detail, there are four key modules in the framework of UrbanWorld. Firstly, UrbanWorld automatically generates untextured 3D layouts with the above-mentioned input data and conducts detailed asset processing via Blenderhttps://www.blender.org/. Then, UrbanWorld adopts a fine-tuned urban-specific multimodal large language model (called Urban MLLM) to effectively plan and design urban scenes following user instructions, generating detailed textual descriptions of urban elements. Next, UrbanWorld integrates a 3D asset renderer based on texture diffusion and refinement, flexibly controlled by textual and visual conditions. Finally, to further optimize the visual appearance, it utilizes Urban MLLM to scrutinize the crafted 3D urban environment, generating detailed suggestions for refinement and activate an additional iteration of rendering.
Our framework is highly flexible to support generating two typical types of 3D urban environments. On the one hand, UrbanWorld can generate a highly accurate replica of the real urban environment with real street-view imagery as the generation condition, which holds significant potential for studies of urban planning and geographic information systems. On the other hand, it can also generate fully customized urban environments with textual descriptions as the generation condition. This capability is valuable for simulating and exploring hypothetical urban scenarios, especially in areas such as virtual city design, gaming, and emergency response planning. In the experimental part, we begin by conducting a comprehensive quantitative evaluation using five visual metrics, validating the state-of-the-art realism of UrbanWorld’s generated environments. We then showcase diverse generation results from various textual and image prompts, highlighting the superior controllability of UrbanWorld. Finally, we emphasize the interactive nature of the generated 3D urban environments by demonstrating agent perception and navigation within them. We contribute UrbanWorld as an open platform to support the creation and manipulation of more advanced 3D urban environments, facilitating the advancement of broad AI communities.
@@
  • We provide UrbanWorld as an open-source platform to develop 3D urban environments, which can support a wide range of research including embodied intelligence and AI agents, further laying a foundation for the advancement of AGI.
@@ 2 Related Works
Table 1: Comparison between existing works for 3D city generation and UrbanWorld from four aspects: text-controllable, image-controllable, whether new assets can be created, and interactive.
Type Method Text-controllable Image-controllable Creating new assets Interactive
Neural Rendering SceneDreamer [13] $\times$ $\checkmark$ $\checkmark$ $\times$
PersistentNature [14] $\times$ $\checkmark$ $\checkmark$ $\times$
Infinicity [6] $\times$ $\checkmark$ $\checkmark$ $\times$
CityDreamer [7] $\times$ $\times$ $\checkmark$ $\times$
Diffusion CityGen [8] $\times$ $\times$ $\checkmark$ $\times$
3D Modeling Software MetaUrban [15] $\times$ $\times$ $\times$ $\checkmark$
SceneCraft [16] $\checkmark$ $\times$ $\times$ $\checkmark$
CityCraft [17] $\checkmark$ $\times$ $\times$ $\checkmark$
Comprehensive UrbanWorld $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
3D urban scene generation aims to create realistic 3D urban environments with sophisticated urban planning and visual element design, usually requiring high human efforts such as complex asset modeling, texture mapping, and scene composition. With the advancement of deep learning techniques, recently there are three lines of work trying to achieve this in an automated way, including neural rendering-based methods [6, 7, 13], diffusion-based methods [8, 18, 19] and 3D modeling software-based methods [20, 16, 15]. An overview of the method comparison is shown in Table 1. Neural rendering-based methods implicitly represent the urban scene and perform the volumetric rendering for the neural fields. For example, CityDreamer [7] first separates the scene into buildings and backgrounds then introduces different types of neural fields for asset rendering. These methods can produce a high-quality visual appearance while potentially losing geometric fidelity. Diffusion-based methods utilize diffusion models to generate city layouts or urban scenes. CityGen [8] provides an end-to-end pipeline to create diverse 3D city layouts with Stable Diffusion. These methods are creative in generating scene images or videos, but hard to obtain embodied 3D environments, limiting the practical usages. Recently, some professional software script-based methods have been proposed, trying to develop an automatic agentic workflow using LLMs to control the professional software for scene creation. CityCraft [17] adopts LLMs in designing and organizing 3D urban environments from off-the-shelf asset libraries. Such approaches are effective but only create urban environments with existing 3D models, unable to flexibly create new assets when necessary. By comparison, our model can freely create new 3D urban assets in a highly controllable way, allowing for crafting diverse urban environments.
@@
A persistent objective in AI research has been to develop machine agents capable of engaging with various environments in 3D space like humans. Toward this goal, researchers have been devoted to building various interactive world simulators in the format of videos [21] or embodied environment [22]. Existing world simulation environments and platforms are mostly for indoor scenes [23, 24, 25, 26]. Differently, Threedworld [27] paid attention to creating outdoor environments by retrieving and compositing objects from an existing asset library. When it comes to urban scenes which is the most important open environment, existing works mostly focus on the generative world model for autonomous driving capable of learning scene dynamics and understanding the geometry of the physical world [10, 11, 28]. However, these models can only generate new scenes in the format of videos, hard to provide an embodied and interactive urban environment for real use. UGI [4] conceptualized some relevant ideas toward urban world simulation, proposing to develop an embodied urban environment for agent development, but still lacks practical implementation. More recently, MetaUrban [15] developed a simulation platform for embodied agents in the urban environment, while the provided environment is limited to a fixed style without controllability. To address these challenges, we propose UrbanWorld, which is expected to facilitate the construction of diverse embodied urban environments with controllable and refined visual appearance, supporting agent development or simulation in various urban scenes.
@@ 3 Methodology
Figure 1: Illustration of the whole framework of UrbanWorld, including four key components: (A) Flexible 3D urban layout generation; (B) Urban MLLM-empowered scene design; (C) Diffusion-based urban asset texture rendering; (D) MLLM-assisted scene refinement.
@@
There are three main challenges to solve for building a real “urban world model”: efficient interactive environment construction, elaborate urban scene planning and high-quality texture rendering. Towards these objectives, UrbanWorld follows a novel “map-design-render-refine” generation pipeline, accomplished by the collaboration of a specialized urban MLLM and an urban asset rendering module. In detail, there are four key stages in UrbanWorld: (1) Flexible 3D urban layout generation, achieving automatic 2.5D-to-3D mapping based on various urban layout metadata such as globally open-accessible OSM data and urban layouts with semantic and height maps, which can address the first challenge. (2) Urban MLLM-empowered scene design, which exploits the superior urban environment understanding ability of a fine-tuned urban MLLM to draft realistic urban scenes emulating human designers for addressing the second challenge. (3) Controllable diffusion-based urban asset texture renderer, achieving flexible urban asset rendering based on 3D diffusion supporting both textual and visual prompts. (4) MLLM-assisted urban scene refinement, exploiting the urban MLLM to conduct reflection to refine the generated urban environment, mimicking the iterative revision in the standard operation process of human designers. The last two components contribute to high-fidelity textures of 3D assets, effectively tackling the third challenge. The overview framework of UrbanWorld is illustrated in Figure 1.
3.1 Flexible 3D Urban Layout Generation
@@
UrbanWorld supports any type of 2.5D layout data as input to efficiently build the untextured 3D urban environment, such as easily accessible OSM data and AI-generated urban layout data. These 2.5D layout data record the metadata of an urban area including the geographic locations and attributes of diverse urban elements such as roads, buildings, vegetation, forests, and water. All urban assets will be automatically separated as independent objects for subsequent element-wise rendering. In this step, UrbanWorld also records the object center location for further reorganization of assets, making it match the real urban layout.
3.2 Urban MLLM-empowered Scene Design
@@
Aiming to effectively craft customized urban environments, UrbanWorld integrates an advanced MLLM fine-tuned on extensive real-world urban imagery data, named Urban MLLM. In detail, we collected approximately 300K urban street-view images from Google Maps and used GPT-4 to generate associated textual descriptions. These descriptions were then manually reviewed and low-quality data were filtered out. Subsequently, we fine-tuned an advanced open-source MLLM, VILA-1.5 [29] with the curated dataset to improve its understanding of urban environments. In the workflow of UrbanWorld, Urban MLLM is utilized to act as a human-like designer, which automatically drafts high-quality and detailed urban scene descriptions, ensuring the crafted urban environment are visually coherent and instruction-following. Specifically, taking a simple textual instruction (e.g., “a teaching area in the university”) from users as input, UrbanWorld calls Urban MLLM with carefully designed prompts and returns diverse detailed descriptions about visual appearance and materials for each asset. The produced asset descriptions will be used to control the condition of the later texture rendering process.
@@ 3.3 Controllable Diffusion-based Urban Asset Texture Rendering
Rendering a large-scale urban scene is challenging due to the existence of complex elements and relations, whereby the scene-level rendering will inevitably result in texture mismatching and low-resolution issues. Therefore, we adopt the element-wise rendering strategy to ensure the rendering quality. Simultaneously, in order to speed up the rendering process, we merge some urban element types and finally define four main categories: buildings, roads and paths, forest and vegetation, and water. We implement the rendering with a controllable diffusion-based method consisting of two stages: UV texture generation and texture refinement as shown in Figure 2.
After the 3D environment generation detailed in Section 3.1, we have obtained the untextured 3D mesh of an urban region $S$ and contained assets $\{S_{1},S_{2},...,S_{i}\}$. For each type of asset, UrbanWorld takes the associated textual description $t_{i}$ from Urban MLLM or a reference image $r_{i}$ as prompts to control the generation.
We first set a series of camera views $v_{i}$ = $\{v_{i}^{k}\}_{k=1}^{N}$ to capture multi-view appearances of the asset $S_{i}$. Next, we utilize the depth-aware ControlNet [30] to control a 2D diffusion model $F$ to generate an image $I_{i}$ showing the visual appearance of $S_{i}$ on different views, controlled by the condition $c_{i}\in\{t_{i},r_{i},(t_{i},r_{i})\}$:
@@
where $z$ is the latent embedding for the diffusion process, the depth map from different views $d_{i}$ is obtained from the rendering process $d_{i}=P(S_{i},v_{i})$. Then we crop $I_{i}$ into patches where each represents a unique view $\{I_{i}^{k}\}_{k=1}^{N}$ of the rendered asset. Then we conduct a reverse process $P^{-1}$ of rendering to back-project $I_{i}$ into the UV texture space:
$U_{i}^{k}=P^{-1}(v_{i}^{k};I_{i}^{k};S_{i}),$ (2)
Subsequently, we merge the texture maps from different views into a single texture map $U_{i}$:
@@
Up to now, we have obtained the preliminary texture map for the asset $S_{i}$, while in practice we found that there are still some untextured areas on the object due to the discrete sampling of camera views, especially for assets with many faces. Inspired by the inpainting capability of diffusion models, we introduce an additional UV texture inpainting process to get more complete and natural texture. However, such inpainting can not be directly achieved with general diffusion-based inpainting since the inpainting should be restricted to follow the position relation in the UV texture space. Therefore, we introduce a position map-guided building UV inpainting process, inspired by the inpainting process of general 3D objects  [31].
@@
To be specific, we add a UV position map encoder $E_{V}(\cdot)$ to encode the position map $V_{i}\in\mathbb{R}^{H\times W\times 3}$, indicating the adjacency relation of the UV texture fragments, where $E_{V}$ is set to have the same architecture of the image encoder in the diffusion model. Then we curate a set of paired-up UV position maps and UV texture maps for urban assets with complex surfaces and train the position encoder following the pipeline of ControlNet [30]. With the control of UV position maps, it’s expected to achieve accurate and natural inpainting for UV texture maps. Denote $U_{i}^{*}$ as the inpainted UV texture map, the UV inpainting process is formulated as follows:
$U_{i}^{*}=F(c_{i};U_{i};E_{V}(V_{i})).$ (4)
With the above texture generation and completion, UrbanWorld can produce coherent and high-fidelity textures for various urban elements. For better visual aesthetic of the rendering, we further conduct image upscaling with ControlNet-tile to enhance the structure sharpness and realism of the texture map, contributing to more detailed and realistic appearances of urban assets.